# HiCache: Training-free Acceleration of Diffusion Models via Hermite Polynomial-based Feature Caching

### v1

@@ Abstract
Diffusion models have achieved remarkable success in content generation but suffer from prohibitive computational costs due to iterative sampling. While recent feature caching methods tend to accelerate inference through temporal extrapolation, these methods still suffer from server quality loss due to the failure in modeling the complex dynamics of feature evolution. To solve this problem, this paper presents HiCache, a training-free acceleration framework that fundamentally improves feature prediction by aligning mathematical tools with empirical properties. Our key insight is that feature derivative approximations in Diffusion Transformers exhibit multivariate Gaussian characteristics, motivating the use of Hermite polynomials—the potentially theoretically optimal basis for Gaussian-correlated processes. Besides, We further introduce a dual-scaling mechanism that ensures numerical stability while preserving predictive accuracy. Extensive experiments demonstrate HiCache’s superiority: achieving 6.24× speedup on FLUX.1-dev while exceeding baseline quality, maintaining strong performance across text-to-image, video generation, and super-resolution tasks. Core implementation is provided in the appendix, with complete code to be released upon acceptance.
@@
Figure 1: Visual demonstration of HiCache using the FLUX.1-dev model, with a $\boldsymbol{6.24\times}$ speedup ($N_{\text{interval}}=7$).
1. Introduction
@@
Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021; Dhariwal and Nichol 2021), especially when implemented with the scalable Transformer architecture (DiT) (Peebles and Xie 2023), have set a new standard for high-fidelity content generation (Chen et al. 2023; Rombach et al. 2022; Karras et al. 2022; Saharia et al. 2022; Balaji et al. 2022). This success, however, is tempered by the substantial computational overhead inherent to the diffusion process. The iterative sampling, requiring hundreds of sequential forward passes through large models, results in high latency that critically hinders practical deployment. Consequently, developing efficient acceleration methods is a key research objective.
Among various acceleration strategies, feature caching has emerged as a particularly effective and training-free approach. Motivated by the observation that diffusion models exhibit similar features in the adjacent timesteps, the original feature caching seeks to directly reuse features in the adjacent timesteps, which leads to significant generation quality loss when the acceleration ratios increase (Selvaraju et al. 2024; Ma et al. 2023; Zou et al. 2024). Recently, the paradigm has advanced to a “cache-then-forecast” mechanism as TaylorSeer (Liu et al. 2025), where Taylor series expansion is used to extrapolate features in the future timesteps, significantly reducing the error from feature caching. TaylorSeer significantly reduces the error introduced by feature reusing, but still faces fundamental theoretical limitations: The standard polynomial basis of a Taylor series is suboptimal for modeling the complex, non-monotonic trajectories of feature evolution in DMs.
Figure 2 visualize the features of the original diffusion models (i.e., ground-truth) and the predictive features from TaylorSeer. It is observed that Taylor-based extrapolation exhibits significant deviations from ground truth trajectories, particularly at turning points where its monotonic nature fails to capture the underlying dynamics. This mismatch between the Taylor-based predictive model and the underlying feature distribution establishes a bottleneck for achieving much higher acceleration ratios without significant quality degradation. Besides, TaylorSeer is designed to employ high-order Taylor expansion to make use of more historical information, making it sensitive to the noise in the feature trajectory, further limiting its effectiveness.
@@
Figure 2: Trajectory prediction comparison of Taylor and HiCache methods on FLUX. “Full trajectory” indicates the feature trajectory from the original Flux (Esser et al. 2024) in the 14th and 28th layer. The y-axis denotes the most principal component of the features in diffusion models. HiCache demonstrates more stable fitting at trajectory turning points, while Taylor expansion tends to over-extrapolate and deviate from original trajectories.
@@
To overcome this limitation, in this paper, we propose HiCache (Hermite Polynomial-based Feature Cache), a framework that improves the feature forecasting by aligning the predictive basis with the intrinsic statistical properties of the feature dynamics. Our approach is based on a key empirical observation: the derivative approximations of feature vectors in DiTs consistently exhibit a multivariate Gaussian distribution (details in Section 3.3.1). Motivated by this finding, HiCache replaces the standard polynomial basis with scaled Hermite polynomials. According to approximation theory (Szegö 1975), Hermite polynomials are the potentially optimal orthogonal basis for representing Gaussian-correlated processes, making them theoretically better suited for this task. Besides, to address the numerical challenges of Hermite polynomials at large extrapolation steps, we further introduce a dual-scaling mechanism that simultaneously constrains predictions within the stable oscillatory regime and suppresses exponential coefficient growth in high-order terms through a single hyperparameter. This principled design enables both accurate and numerically stable feature prediction. As shown in Figure 2, the feature trajectory predicted by HiCache closely matches the original FLUX, demonstrating its great effectiveness.
[…]
The main contributions of this work are as follows:
  • We systematically validate the multivariate Gaussian nature of feature derivative approximations in Diffusion Transformers, offering a new statistical foundation for designing more efficient feature caching methods.
  • We propose HiCache, which introduces Hermite polynomials into the feature caching of diffusion models, and propose a dual-scaling mechanism to simultaneously constrain predictions within the stable oscillatory regime and suppress exponential coefficient growth in high-order terms, achieving robust numerical stability.
  • We conduct extensive experiments on four diffusion models and generative tasks, demonstrating HiCache’s universal superiority and broad applicability.
2. Related Work
Sampling Step Reduction
@@ 2. Related Work
aims to reduce the total number of denoising iterations. DDIM (Song, Meng, and Ermon 2021) introduced deterministic sampling with larger steps, while DPM-Solver (Lu et al. 2022) formalized this as an ODE with high-order numerical methods. Recent advances include knowledge distillation (Meng et al. 2023) and Consistency Models (Song et al. 2023) for few-step generation. Other notable approaches include cascaded diffusion models (Ho et al. 2022a) and improved sampling techniques (Karras et al. 2022). Text-to-image diffusion models have seen remarkable progress (Saharia et al. 2022; Balaji et al. 2022; Ramesh et al. 2022; Podell et al. 2023; Zhang, Rao, and Agrawala 2023), while video generation (Singer et al. 2022; Ho et al. 2022b) presents additional temporal challenges.
Denoising Network Compression
@@
focuses on reducing per-step computational cost. Model compression methods include neural network pruning (Fang et al. 2023) and quantization (Li et al. 2023) have applied the traditional model compression methods to diffusion models. Besides, token merging (Bolya and Hoffman 2023) has been introduced to reduce the number of tokens computed in each timestep. However, these methods tend to lead clear loss in generation quality, and sometimes rely on retraining the diffusion model.
Feature Caching
@@
has been recently proposed to offer a training-free alternative. Motivated by the great similarity of features in the adjacent timesteps, these methods aim to skip the computation in some timesteps based on the historical features, which can be roughly divided into two paradigms: (1) Cache-then-Reuse directly reuses features from adjacent timesteps (DeepCache (Ma et al. 2023), FORA (Selvaraju et al. 2024), ToCa (Zou et al. 2024)), but suffers quality degradation at large intervals; (2) Cache-then-Forecast predicts future features using historical data. TaylorSeer (Liu et al. 2025) pioneered this approach with Taylor series expansion, significantly improving quality at high acceleration ratios. However, its standard polynomial basis remains suboptimal for complex feature dynamics, motivating our Hermite polynomial-based approach.
3. Methodology
3.1 Preliminaries: Diffusion Transformer
The Diffusion Transformer (DiT) architecture is central to many state-of-the-art diffusion models. Building upon the self-attention mechanism (Vaswani et al. 2017), it operates by progressively denoising a noisy input $x_{t}$ through a series of $L$ stacked blocks. Within each block, an input feature map $\mathcal{F}_{\text{in}}$ is updated via sequential residual corrections:
$\mathcal{F}^{\prime}$ $=\mathcal{F}_{\text{in}}+\text{MHSA}(\text{AdaLN}(\mathcal{F}_{\text{in}},t))$ (1)
$\mathcal{F}_{\text{out}}$ $=\mathcal{F}^{\prime}+\text{MLP}(\text{AdaLN}(\mathcal{F}^{\prime},t))$ (2)
where AdaLN denotes adaptive layer normalization conditioned on timestep $t$. This temporal conditioning, combined with nonlinearities (e.g., Softmax and GeLU), causes feature trajectories to exhibit complex, non-monotonic dynamics—a significant challenge for acceleration techniques relying on smooth extrapolation. The importance of such architectural choices in generative models has been widely studied (Karras et al. 2022; Dhariwal and Nichol 2021).
@@ 3.2 Taylor Expansion-based Feature Prediction
This limitation of the monomial basis $\{(-k)^{i}\}_{i=1}^{m}$ motivates our search for a more suitable basis function.
@@ 3.3 HiCache: Hermite-based Feature Caching
In this paper, we propose to leverage the Gaussian nature of feature dynamics for optimal basis selection.
Observation 1 (Gaussianity of Feature Derivatives).
Let $\{\mathcal{F}_{t}\}_{t=1}^{T}$ be the feature trajectory in a Diffusion Transformer. The finite difference approximations $\{\Delta^{k}\mathcal{F}_{t}\}$ follow a multivariate Gaussian distribution:
$\Delta^{k}\mathcal{F}_{t}\sim\mathcal{N}(\mu_{k},\Sigma_{k})$ (6)
Validated via Energy statistical tests with p-value = 1.0 across all configurations (see Section 5.1).
@@
By the Karhunen-Loève theorem, Hermite polynomials $\{H_{n}\}$ form the optimal orthogonal basis for Gaussian-correlated processes (Narcowich, Ward, and Wendland 2005), implying superior approximation properties over monomials.
@@ 3.3.2 Hermite Polynomials as Basis Functions
Given the Gaussian characteristics, we adopt Hermite polynomials as our basis functions. The physicist’s Hermite polynomials $H_{n}(x)$ (Abramowitz and Stegun 1964) are defined via the Rodrigues formula:
$H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$ (7)
and satisfy the recurrence relation:
$H_{n+1}(x)=2xH_{n}(x)-2nH_{n-1}(x)$ (8)
Unlike monotonically growing Taylor basis functions, Hermite polynomials exhibit inherent oscillatory behavior (as shown in Figure 3). This oscillatory nature provides implicit regularization (van der Westhuizen and Lasenby 2018), enabling more accurate capture of non-monotonic dynamics in neural network optimization. However, Hermite polynomials face numerical challenges: they grow exponentially for large $|x|$ and have rapidly increasing coefficients due to the factor of 2 in the recurrence. To address this, we introduce scaled Hermite polynomials:
@@
Figure 3: Systematic comparison of Taylor and Hermite basis functions Hermite’s oscillatory behavior (e.g., $H_{2}(x)$ negative offset) captures non-monotonic evolution better than monotonic Taylor growth. (Y-axis: symmetric log scale).
@@ 3.3.3 HiCache Algorithm
Given cached derivative approximations $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$, the optimal predictor under Gaussian feature dynamics is:
@@
where $k\in\{1,\ldots,N_{\text{interval}}-1\}$ is the prediction horizon and $\tilde{H}_{i}$ is the scaled Hermite polynomial from Definition 2.
Proof Intuition. Under the empirically validated Gaussian assumption (Observation 1), Corollary 1 suggests that Hermite polynomials provide a theoretically motivated basis for feature prediction. The factorial normalization ensures proper scaling of higher-order terms. See Appendix A.6 for detailed theoretical analysis and supporting arguments.
@@
1:for timestep $t=T$ to 11 do
2:  if $t\mod N_{\text{interval}}=0$ then
3:   Compute $\mathcal{F}_{t}=\Phi_{t}(x_{t})$ {Full model forward}
4:   Update cache: $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$ via Def. 1
5:  else
6:   $k\leftarrow t\mod N_{\text{interval}}$
7:   Predict $\hat{\mathcal{F}}_{t}$ using Eq. (10)
8:  end if
9:end for
Computational Complexity. HiCache reduces computational cost by factor $(N_{\text{interval}}-1)/N_{\text{interval}}$ while maintaining $O(N_{\text{order}})$ prediction overhead per skipped step.
Proposition 3 (Error Bounds for HiCache).
The total prediction error of HiCache decomposes as:
with the bound:
$\|\mathbf{E}_{\text{total}}\|\leq\underbrace{O\left(\frac{(\sigma\sqrt{2}|\Delta s|)^{N+1}}{\sqrt{(N+1)!}}\right)}_{\text{Hermite truncation}}+\underbrace{O(\Delta t_{\text{hist}}\sqrt{N})}_{\text{Finite difference}}+\underbrace{O(\epsilon_{\text{machine}})}_{\text{Floating point}}$ (12)
Corollary 2 (Superiority over Taylor Expansion).
HiCache achieves:
  1. 1.
    Faster convergence: $O(1/\sqrt{(N+1)!})$ vs. $O(1/(N+1)!)$
  2. 2.
    Exponential suppression via $\sigma^{N+1}$
  3. 3.
    Automatic adaptation to Gaussian feature statistics
Proof. See Appendix A.3 for complete derivation.
Figure 4: An overview of HiCache compared with TaylorSeer. TaylorSeer (orange) predicts the feature with a power basis. HiCache (green) uses the same functional form but replaces the basis with Hermite polynomials. Dashed curves show predicted trajectories against ground truth, highlighting error accumulation at skipped steps.
4. Experiments
4.1 Experiment Settings
We conduct comprehensive experiments across four representative generative tasks using state-of-the-art diffusion models: text-to-image generation with FLUX.1-dev (Esser et al. 2024), text-to-video generation with HunyuanVideo (Kong et al. 2024), class-conditional image generation with DiT-XL/2 (Peebles and Xie 2023), and image super-resolution with a modified Inf-DiT (Yang et al. 2024). Each task employs standard evaluation protocols and widely-adopted benchmarks to ensure fair comparison with existing methods. More detailed information is provided in Appendix A.4.
4.2 Text-to-Image Generation
Quantitative Analysis.
As shown in Table 1, HiCache demonstrates a clear quantitative advantage. We evaluate using ImageReward (Xu et al. 2023) as the primary metric for perceptual quality assessment. At a  5.55x acceleration, HiCache not only surpasses its main competitor, TaylorSeer, but also slightly exceeds the unaccelerated baseline, suggesting near-lossless performance. This advantage becomes more pronounced at higher accelerations ( 6.24x), where HiCache maintains a robust score while competitors’ performance degrades more significantly.
Table 1: Quantitative comparison on the text-to-image generation task with the FLUX.1-dev model.
Method Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP $\uparrow$
@@ 4.2 Text-to-Image Generation
Numbers in parentheses show the deviation from the DDIM-50 baseline (Image Reward=0.9872). PSNR and LPIPS are compared against the ideal values of $\infty$ and 0, respectively. Among the compared methods, only ToCa does not utilize efficient attention.
Qualitative Analysis.
As illustrated in Figure 5, HiCache consistently produces high-fidelity results, avoiding common artifacts like facial distortions, conceptual errors, or text inaccuracies that affect competing methods. Furthermore, Figure 6 highlights HiCache’s superior ability to preserve high-frequency details, such as intricate textures and subtle reflections, which are often lost in other accelerated outputs.
Figure 5: Qualitative comparison on the text-to-image task across nine diverse prompts.
Figure 6: Visualization of high-frequency detail preservation.
Robustness Under Increasing Acceleration.
Figure 7 assesses stability under increasing acceleration factors (from 1.00x to 9.00x). TaylorSeer’s output shows significant style drift and structural degradation, consistent with the error accumulation discussed in Sec. 5.1. In contrast, HiCache maintains high stability and aesthetic consistency, demonstrating its robustness against such errors even at the highest acceleration levels.
Figure 7: Comparison between TaylorSeer (second-best method) and HiCache in different acceleration ratios.
4.3 Text-to-Video Generation
Quantitative Analysis.
We further evaluate HiCache on text-to-video generation using the HunyuanVideo model (Kong et al. 2024). As detailed in Table 3, which benchmarks performance with VBench (Huang et al. 2024), HiCache consistently outperforms competitors. Notably, when comparing higher-order configurations (e.g., $N=7,O=2$), HiCache demonstrates even more significant advantages over TaylorSeer, which aligns with our theoretical analysis in Section 3.3.2 that Hermite polynomials’ oscillatory behavior better captures complex feature dynamics at higher orders, while Taylor expansion’s monotonic nature leads to increasing over-extrapolation.
Method Efficient Attention Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ FID $\downarrow$ sFID $\downarrow$ Inception Score $\uparrow$

### v2

# HiCache: A Plug-in Scaled-Hermite Upgrade for Taylor-Style Cache-then-Forecast Diffusion Acceleration

@@ Abstract
Diffusion models have achieved remarkable success in content generation but suffer from prohibitive computational costs due to iterative sampling. While recent feature caching methods tend to accelerate inference through temporal extrapolation, these methods still suffer from severe quality loss due to the failure in modeling the complex dynamics of feature evolution. To solve this problem, this paper presents HiCache (Hermite Polynomial-based Feature Cache), a training-free acceleration framework that fundamentally improves feature prediction by aligning mathematical tools with empirical properties. Our key insight is that feature derivative approximations in Diffusion Transformers exhibit multivariate Gaussian characteristics, motivating the use of Hermite polynomials—the potentially theoretically optimal basis for Gaussian-correlated processes. Besides, we introduce a dual-scaling mechanism that ensures numerical stability while preserving predictive accuracy, which is also effective when applied standalone to TaylorSeer. Extensive experiments demonstrate HiCache’s superiority: achieving $5.55\times$ speedup on FLUX.1-dev while exceeding baseline quality, maintaining strong performance across text-to-image, video generation, and super-resolution tasks. Moreover, HiCache can be naturally added to the previous caching methods to enhance their performance, e.g., improving ClusCa from $0.9480$ to $0.9840$ in terms of image rewards. Code: https://github.com/fenglang918/HiCache.
footnotetext: Emails: fenglang918@foxmail.com; zhanglinfeng@sjtu.edu.cn.
@@
Figure 1: Teaser of HiCache using the FLUX.1-dev model, with a $\bm{6.24\times}$ FLOPs speedup .
1 Introduction
@@
Diffusion Models (DMs)  (Ho et al., 2020; Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021), especially when implemented with the scalable Transformer architecture (DiT)  (Peebles and Xie, 2023), have set a new standard for high-fidelity content generation  (Chen and others, 2023; Rombach et al., 2022; Karras et al., 2022; Saharia and others, 2022; Balaji and others, 2022). This success, however, is tempered by the substantial computational overhead inherent to the diffusion process. The iterative sampling, requiring hundreds of sequential forward passes through large models, results in high latency that critically hinders practical deployment, especially on diffusion transformers. Consequently, developing acceleration methods for diffusion models has become a hot research topic.
@@
Figure 2: Trajectory prediction comparison of Taylor and HiCache methods on FLUX. “Full trajectory” indicates the feature trajectory from the original Flux (Esser and others, 2024) in the 14th and 28th layer. The y-axis denotes the most principal component of the features in diffusion models.
Among various acceleration strategies, feature caching has emerged as a particularly effective and training-free approach. Motivated by the observation that diffusion models exhibit similar features in the adjacent timesteps, the original feature caching seeks to directly reuse features in the adjacent timesteps (e.g., DeepCache (Ma and others, 2023), FORA (Selvaraju and others, 2024), ToCa (Zou et al., 2025), ClusCa (Zheng et al., 2025)), which leads to significant generation quality loss when the acceleration ratios increase. Recently, the paradigm has advanced to a “cache-then-forecast” mechanism as TaylorSeer (Liu et al., 2025a), where Taylor series expansion is used to extrapolate features in the future timesteps, significantly reducing the error from feature caching. TaylorSeer significantly reduces the error introduced by feature reusing, but still faces fundamental theoretical limitations: The standard polynomial basis of a Taylor series is suboptimal for modeling the complex, non-monotonic trajectories of feature evolution in DMs.
Figure 2 visualize the features of the original diffusion models (i.e., ground-truth) and the predictive features from TaylorSeer. It is observed that Taylor-based extrapolation exhibits significant deviations from ground truth trajectories, particularly at turning points where its monotonic nature fails to capture the underlying dynamics. These deviations indicate a modeling mismatch: the monomial Taylor basis poorly fits the non-monotonic, turning-point behavior of feature trajectories, causing approximation error to grow rapidly with prediction horizon and order (Proposition 1); this, in turn, caps the attainable acceleration before quality degradation becomes evident.
@@
To overcome this limitation, in this paper, we propose HiCache (Hermite Polynomial-based Feature Cache), a framework that improves the feature forecasting by aligning the predictive basis with the intrinsic statistical properties of the feature dynamics. Our approach is based on a key empirical observation: the derivative approximations of feature vectors in DiTs consistently exhibit a multivariate Gaussian distribution (details in Section 3.3.1). Motivated by this finding, HiCache replaces the standard polynomial basis with scaled Hermite polynomials. According to approximation theory (Szegö, 1975), Hermite polynomials are the potentially optimal orthogonal basis for representing Gaussian-correlated processes, making them theoretically better suited for this task. Besides, to address the numerical challenges of Hermite polynomials at large extrapolation steps, we further introduce a dual-scaling mechanism that simultaneously constrains predictions within the stable oscillatory regime and suppresses exponential coefficient growth in high-order terms through a single hyperparameter. This principled design enables both accurate and numerically stable feature prediction. As shown in Figure 2, the feature trajectory predicted by HiCache closely matches the original FLUX, demonstrating its effectiveness. Importantly, HiCache naturally serves as a drop-in replacement for Taylor-based predictors in existing cache-then-forecast frameworks, requiring only the substitution of the polynomial basis while preserving the same predictor form and computational structure with negligible overhead.
Extensive experiments demonstrate the effectiveness of HiCache in class-to-image generation with DiT, text-to-image generation with FLUX.1-dev, image-to-image generation with Inf-DiT for image super-resolution, and text-to-video generation with HunyuanVideo (Kong and others, 2024), showing improvements over previous methods. For instance, on FLUX.1-dev, HiCache achieves $5.55\times$ acceleration with $4.25\%$ improvement on ImageReward over the second-best method. This superiority extends to its use as a component upgrade: replacing ClusCa’s Taylor predictor with our Hermite basis yields ImageReward improvements from $0.9480$ to $0.9840$ at $\mathcal{N}=7$.
In summary, our main contributions are as follows.
  • Hermite Polynomial-based Feature Caching. We propose a training–free feature extrapolation/cache predictor grounded in the empirically Gaussian finite–difference statistics of diffusion features; instantiated via KL–optimal scaled Hermite polynomials under Gaussian correlation, it better captures non–monotonic trajectories while preserving the Taylor–style form and compute.
  • Dual-scaling mechanism. We introduce a single-hyperparameter dual scaling (input contraction + coefficient suppression) that stabilizes predictions and suppresses high-order growth without accuracy loss; it also works standalone for TaylorSeer.
  • Plug-and-play generality. HiCache is model-agnostic and training-free; by only swapping the polynomial basis, it drops into any Taylor-style cache-then-forecast pipeline (e.g., TaylorSeer, ClusCa) with unchanged predictor form and negligible compute overhead—only a few per-step scalar basis-function evaluations and no extra matrix multiplications.
  • Broad empirical gains. Across text-to-image, text-to-video, class-conditional, and super-resolution, we achieve equal-or-higher speedups with better quality: e.g., on FLUX.1-dev, $5.55\times$ speedup with ImageReward improving from $0.9872$ to $0.9979$; a zero-extra FLOPs basis swap lifts ClusCa from $0.9480$ to $0.9840$.
2 Related Work
@@
Sampling Step Reduction aims to reduce denoising iterations. DDIM (Song et al., 2021) introduced deterministic sampling with large steps, while DPM-Solver (Lu and others, 2022) formalized as an ODE with high-order numerical methods. Advances include knowledge distillation (Meng and others, 2023) and Consistency Models (Song and others, 2023) for few-step generation. Other approaches include cascaded diffusion models (Ho and others, 2022a) and improved sampling techniques (Karras et al., 2022). Text-to-image diffusion models show progress (Saharia and others, 2022; Balaji and others, 2022; Ramesh and others, 2022; Podell et al., 2023; Zhang et al., 2023), while video generation (Singer and others, 2022; Ho and others, 2022b) presents additional challenges.
@@
Denoising Network Compression focuses on reducing per-step computational cost. Model compression methods include neural network pruning (Fang and others, 2023) and quantization (Li and others, 2023) have applied the traditional model compression methods to diffusion models. Besides, token merging (Bolya and Hoffman, 2023) has been introduced to reduce the number of tokens computed in each timestep. However, these methods tend to lead clear loss in generation quality, and sometimes rely on retraining the diffusion model.
@@
Feature Caching has been recently proposed to offer a training-free alternative. Motivated by the great similarity of features in the adjacent timesteps, these methods aim to skip the computation in some timesteps based on the historical features, which can be roughly divided into two paradigms: (1) Cache-then-Reuse directly reuses features from adjacent timesteps (DeepCache (Ma and others, 2023), FORA (Selvaraju and others, 2024), ToCa (Zou et al., 2025), L2C (Ma et al., 2024)), but suffers quality degradation at large intervals; (2) Cache-then-Forecast predicts future features using historical data. TaylorSeer (Liu et al., 2025a) pioneered this approach with Taylor series expansion, significantly improving quality at high acceleration ratios. However, its standard polynomial basis remains suboptimal for complex feature dynamics, motivating our Hermite polynomial-based approach.
3 Method
3.1 Preliminaries
Diffusion Transformers (DiT)  (Peebles and Xie, 2023) apply transformer blocks iteratively during denoising, with each block updating features via $\mathcal{F}_{\text{out}}=\mathcal{F}_{\text{in}}+\text{MHSA}(\text{AdaLN}(\mathcal{F}_{\text{in}},t))+\text{MLP}(\text{AdaLN}(\cdot,t))$. The timestep-conditioned normalization and nonlinearities create complex, non-monotonic feature trajectories that challenge existing acceleration methods.
@@ 3.2 Taylor Expansion-based Feature Prediction
This limitation motivates our search for a basis that naturally captures non-monotonic behaviors.
@@ 3.3 HiCache: Hermite-based Feature Caching
Our key insight is that feature differences in diffusion transformers exhibit Gaussian properties (empirically validated in Section 5), which theoretically motivates the use of Hermite polynomials as optimal basis functions.
Proposition 2 (Gaussianity of Feature Differences).
Feature differences $\Delta F_{t}=F(x_{t},t)-F(x_{t-\Delta},t-\Delta)$ in diffusion transformers exhibit approximate Gaussian behavior through: (i) local linearization yielding conditional Gaussianity $\Delta F_{t}\approx\mathcal{N}(\mu_{t},\Sigma_{t})$ for small $\Delta$, and (ii) aggregation effects across network components invoking CLT with Berry-Esseen bound $O(\eta_{t})$. See Appendix G for the complete statement and proof.
@@
When the temporal correlation can be approximated by a Gaussian kernel $K(s,t)=\exp(-(s-t)^{2}/2\tau^{2})$, the Karhunen-Loève expansion yields (scaled) Hermite functions as eigenfunctions. Thus, under such Gaussian kernel correlation, Hermite polynomials form the optimal orthogonal basis in the weighted $L^{2}(\gamma)$ sense. For general Gaussian processes, they remain approximately optimal (see Appendix G, Theorem 2, for precise conditions).
@@ 3.3.1 Hermite Polynomials as Basis Functions
Given the Gaussian nature, we adopt Hermite polynomials $H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$ (Abramowitz and Stegun, 1964) satisfying the recurrence $H_{n+1}(x)=2xH_{n}(x)-2nH_{n-1}(x)$. Unlike monotonically growing Taylor basis functions, Hermite polynomials exhibit oscillatory behavior (as shown in Figure 3(b)). This oscillatory nature provides implicit regularization (van der Westhuizen and Lasenby, 2018), enabling accurate capture of non-monotonic dynamics in neural network optimization.
Unlike monotonic Taylor basis, Hermite polynomials’ oscillatory nature captures turning points. However, they face numerical instability for large arguments, motivating our scaled version:
@@
Remark. The dual-scaling mechanism can also be applied as a standalone enhancement to TaylorSeer, providing measurable gains without changing its basis.
(a) Overview of HiCache compared with TaylorSeer.
(b) Comparison of basis functions.
Figure 3: Method overview and basis function comparison. (a) TaylorSeer (orange) predicts features using a power basis, whereas HiCache (green) keeps the functional form but swaps the basis for scaled Hermite polynomials. (b) Hermite’s oscillatory behavior (e.g., $H_{2}(x)$ negative offset) captures non-monotonic evolution better than monotonic Taylor growth.
@@ 3.3.2 HiCache Algorithm
Given cached derivative approximations $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$, under Gaussian feature dynamics, a theoretically-motivated predictor is given by:
@@
where $k\in\{1,\ldots,N_{\text{interval}}-1\}$ is the prediction horizon and $\tilde{H}_{i}$ is the scaled Hermite polynomial from Definition 2. When coefficients are obtained via weighted least squares, this achieves projection-optimality in $L^{2}(\gamma)$. The direct use of $\Delta^{i}\mathcal{F}_{t}/i!$ provides an efficient approximation.
@@
1: Initialize $t_{\text{last}}\leftarrow T$ {Track last activation step}
2:for timestep $t=T$ to 11 do
3:  if $t\mod N_{\text{interval}}=0$ then
4:   Compute $\mathcal{F}_{t}=\Phi_{t}(x_{t})$ {Full model forward}
5:   Update cache: $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$ via Def. 1 with $\Delta t_{\text{hist}}=N_{\text{interval}}$
6:   $t_{\text{last}}\leftarrow t$ {Update last activation step}
7:  else
8:   $k\leftarrow t_{\text{last}}-t$ {Distance from last activation}
9:   Predict $\hat{\mathcal{F}}_{t}$ using Eq. (5) with $k$
10:  end if
11:end for
HiCache reduces cost by factor $(N_{\text{interval}}-1)/N_{\text{interval}}$ with $O(N_{\text{order}})$ overhead. The total error $\mathbf{E}_{\text{total}}=\mathbf{E}_{\text{truncation}}+\mathbf{E}_{\text{approximation}}+\mathbf{E}_{\text{numerical}}$ satisfies:
$\|\mathbf{E}_{\text{total}}\|\leq O\left(\frac{(\sigma\sqrt{2}|\Delta s|)^{N+1}}{\sqrt{(N+1)!}}\right)+O(\Delta t_{\text{hist}}\sqrt{N})+O(\epsilon_{\text{machine}})$ (6)
Under $\sigma\sqrt{2}|\Delta s|<1$, the Hermite truncation $(\sigma\sqrt{2}|\Delta s|)^{N+1}\allowbreak/\allowbreak\sqrt{(N+1)!}$ can be smaller than Taylor’s $|\Delta s|^{N+1}\allowbreak/\allowbreak(N+1)!$ due to $\sigma^{N+1}$ suppression, yielding conditional superiority (see Appendix D).
Table 1: Comparison on text-to-video generation with HunyuanVideo on VBench.
@@
Table 2: Ablation study with FLUX.1-dev. $\sigma=1.0$ means no contraction factor is applied. Hi-Taylor indicates TaylorSeer upgraded with HiCache.
Method Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
4 Experiments
4.1 Experiment Settings
We conduct comprehensive experiments across four representative generative tasks using state-of-the-art diffusion models: text-to-image generation with FLUX.1-dev (Esser and others, 2024), text-to-video generation with HunyuanVideo (Kong and others, 2024), class-conditional image generation with DiT-XL/2 (Peebles and Xie, 2023), and image super-resolution with a modified Inf-DiT (Yang and others, 2024). Each task employs standard evaluation protocols and widely-adopted benchmarks to ensure fair comparison with existing methods. More detailed information is provided in Appendix E.
4.2 Experiment Results
Text-to-Image Generation. Table 3 shows HiCache maintains superior quality at high acceleration ratios. At $5.55\times$ speedup ($\mathcal{N}=7,\mathcal{O}=2$), its ImageReward ($0.9979$) exceeds both competitors and the unaccelerated baseline ($0.9872$). When integrated with ClusCa by (Zheng et al., 2025), HiCache substantially improves ImageReward: from $0.9480$ to $0.9840$ ($\mathcal{N}=7$) and from $0.8440$ to $0.8860$ ($\mathcal{N}=9$). Figures 4 and 5(a) demonstrate HiCache’s visual advantages. As shown in Figure 5(b), HiCache maintains stability across acceleration factors from $1.00\times$ to $9.00\times$, while other feature caching methods exhibit significant degradation in high acceleration ratios.
Table 3: Quantitative comparison on the text-to-image generation task with the FLUX.1-dev model. We also report Hi-ClusCa (ClusCa + HiCache) to highlight plug-and-play generality.
Method Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$